### v1

# SuperLoc: The Key to Robust LiDAR-Inertial Localization Lies in Predicting Alignment Risks

Abstract
Map-based LiDAR localization, while widely used in autonomous systems, faces significant challenges in degraded environments due to lacking distinct geometric features. This paper introduces SuperLoc, a robust LiDAR localization package that addresses key limitations in existing methods. SuperLoc features a novel predictive alignment risk assessment technique, enabling early detection and mitigation of potential failures before optimization. This approach significantly improves performance in challenging scenarios such as corridors, tunnels, and caves. Unlike existing degeneracy mitigation algorithms that rely on post-optimization analysis and heuristic thresholds, SuperLoc evaluates the localizability of raw sensor measurements. Experimental results demonstrate significant performance improvements over state-of-the-art methods across various degraded environments. Our approach achieves a 54% increase in accuracy and exhibits the highest robustness. To facilitate further research, we release our implementation along with datasets from eight challenging scenarios 111For videos, code and dataset, see superodometry.com/superloc.
I Introduction
Map-based LiDAR localization is a widely used technique in autonomous systems, including autonomous driving, drone inspection, and search and rescue operations. The Iterative Closest Point (ICP) algorithm[1, 2, 3, 4, 5] is the predominant method for estimating a robot’s pose on a 3D prior map. ICP aligns LiDAR scans by matching corresponding points or features, determining relative pose transformations, and enabling the creation of detailed scene maps. While LiDAR-based localization and mapping algorithms[6, 7, 8, 9, 10] have seen success, achieving robust localization still remains challenges. The key challenges are:
Unreliable Performance in Degraded Scenarios: State-of-the-art LiDAR Localization systems[11, 12] can achieve low drift over long distances but often fail in degraded scenarios such as corridors, hallways, tunnels, caves, or planar areas. These environments lack geometric features, leading to insufficient constraints in the optimization process.
Too Late to Prevent Registration Failure: Existing degeneracy mitigation algorithms [13, 14] use solution remapping to update poses along well-constrained directions based on hessian matrix. However, these methods are often too late to prevent imminent failures and suffer from numerical instability because of insufficient constraints during degeneracy.
Motivated by these limitations, we propose SuperLoc, a robust LiDAR localization package by extending our previous odometry system [15] to a map-based localization scenario. The proposed method addresses the above:
1.
Predictive Alignment Risk Assessment: The main contribution of this work is a method that can not only predict the risk of failed alignment between two point clouds but also estimate the confidence of the current scan during the robot operation without suffering from numerical instability in degeneracy.
2.
Active Sensor Fusion: Rather than passively fusing multiple sensors, we estimate observability in the front-end and integrate it into the sensor fusion. This allows us to actively incorporate pose prior regularization from an alternative odometry source at optimal moments. Our active degeneracy mitigation strategy significantly enhances robustness and accuracy, demonstrating a 49.7% improvement over existing methods.
3.
Open-Source Contribution We release our code along with datasets from 8 challenging environments, including ground-truth maps, to push the limits of localization capabilities in extremely challenging scenarios.
We find that the degeneracy problem is not about reducing outliers; rather, it arises from insufficient constraints [16, 17]. Providing additional constraints before failures occur can significantly improve the robustness of localization algorithms. This indicates the importance of predicting alignment risk or degeneracy before, rather than during or after, optimization.
II Related Work
II-A Degeneracy Mitigation in Scan Registration
When robots operate in complex and challenging environments, an ideal LiDAR localization algorithm should be fail-safe and failure-aware to mitigate the effect of ill-conditioned optimization problems. Most existing LiDAR localization methods passively fuse data from all sensor modalities into a single, large optimization problem [18, 19]. This approach assumes that different sensor modalities provide complementary constraints in the optimization process. However, this tightly coupled strategy often requires substantial computational resources. Moreover, it is usually difficult to balance weight between multi modalities. Zhang [13] proposes a solution remapping to only update the pose in well-constrained state directions. However, since the method relies on heuristic thresholds, it is difficult to achieve reliable performance when changing the environment[20]. Similar limitations also exist in [14]. In contrast, our approach can achieve early failure detection and doesn’t need heuristic tuning by evaluating the localizability of raw measurements.
II-B Existing LiDAR Localization Methods
The open-source community offers several excellent LIO SLAM algorithms such as Fast-LIO[12] and LIO-SAM[21], as well as LVIO algorithms like R3LIVE[19] and FAST-LIVO [18]. However, map-based localization algorithms remain relatively scarce, with even fewer performing well in challenging environments. Recent advancements include PALoc[22], which proposes a flexible prior-assisted localization approach using a global factor graph, and HDL_LOC[11], which introduces a portable measurement system for long-term, wide-area localization on prior maps. Additionally, other localization algorithms [23, 24] have demonstrated high-accuracy pose estimation in large-scale environments. Despite these developments, these methods are only tested in standard conditions and prioritize accuracy over robustness.
III Methodology
III-A System Overview
The overview of our system is shown in Fig. 2. The pipeline begins with measurements from LiDAR and IMU. From these inputs, we extract point, plane, and line features. The point-plane correspondence is then used to compute the Principal Component Analysis (PCA), which helps identify the principal and normal directions of the correspondences. These directions are analyzed to estimate observability and obtain confidence values for each state direction. Based on this confidence, we generate an observability scan and actively incorporate pose priors into the sensor fusion system to mitigate degeneracy. The joint optimization process involves multi-metric ICP factor[15], pose priors, and IMU pre-integration[25]. Here we mainly explain the predicted alignment risk, observability estimation, and active sensor fusion module. For other modules, readers can refer to [15] for the detailed structure and implementation.
III-B Predicting Alignment Risk
Detecting alignment risk before odometry failure is crucial for ensuring high accuracy and robustness, especially under extremely challenging environments. Here, we explain how to quantify the alignment risk when registering two point clouds and using point-plane correspondences. We employ a KD-tree to identify point-pairs , where each has an associated normal vector . Our objective is to find a rigid-body transformation, comprising a rotation and a translation , that minimizes the sum of squared distances from each to the plane tangent to at . The alignment error is expressed as:
Then, the point-plane distance transformed by the vector can be rewrite into:
which reveals that points where is perpendicular to , or is perpendicular to , do not contribute to the error . This insight allows us to directly assess whether the current motion, represented by , is constrained by analyzing the orthogonality between these vectors.
For each scan, we can evaluate the rotational and translational stability of correspondences individually to determine if they are constrained by the current motion. This analysis yields the matrix to assess alignment quality[26]:
III-C Observability Estimation
While we can anticipate potential alignment risks, the observability of current state estimation remains uncertain. Most sensor fusion methods [12, 27, 28] currently use covariance matrices to quantify state estimation uncertainty, assuming that optimization adheres to Gaussian distribution even in degenerate cases. However, this assumption may not hold when certain states or landmarks become unobservable. Consequently, the covariance matrix can face numerical instability issues, reducing the reliability of standard Gaussian-based approaches for marginalization and Schur complement[29]. Given that optimization becomes ill-conditioned due to degeneracy, we propose that the covariance matrix should be estimated from the front-end rather than derived from the Hessian matrix in the optimization framework.
Therefore, based on our previous risk prediction, we directly analyze the observability of raw sensor measurements in the front-end. For each correspondence point in the source cloud , we compute six values and project it to the world frame .
where is the observability analysis of points in the world coordinates, , is the initial guess of scan registration, is a planner scalar defined by , and are eigenvalues of PCA for normal computation to represent planarity of geometry[30].
The first three values in indicate the contribution of point to the observability of sensor rotation (roll, pitch, yaw). The last three values represent its contribution to the observability of sensor translations. Based on these values, we can assign labels to each correspondence according to which state direction it contributes most.
To evaluate the observability of whole scan, we statistically analyze the number of observability labels on each state direction as and normalize the results to a [0-1] range as confidence metric:
These equations yield a normalized measure of observability for both translational and rotational components of the state, allowing us to construct a 6x6 diagonal covariance matrix . By examining the distribution of labels across all correspondences in the scan, we can identify which state directions are less constrained and thus require supplementary odometry information. It’s important to note that the equation assesses the ratio of observability labels in a specific direction to the total number of labels. This relative approach is based on the hypothesis that in well-structured environments, observability should be uniformly distributed across all state directions. Consequently, our proposed method eliminates the need for heuristic thresholds.
III-D Active Sensor Fusion
Traditional sensor fusion approaches passively combine all measurements and rely on robust kernel functions to mitigate adverse effects[31]. We propose a more proactive strategy: degeneracy-aware sensor fusion. This actively incorporates additional constraints with proper weighting, particularly in degenerate directions. Our method not only identifies degraded directions but also determines the necessary weight in those directions using the previous proposed matrix.
III-D1 Pose Prior Factor
When degeneracy occurs, we leverage relative poses from alternative odometry sources, denoted as [, ] in local coordinates. These are used to constrain the relative position and rotation estimates obtained from joint optimization. We define the covariance matrix of this prior information as . This regularization term ensures that constraints are equally distributed across state directions to mitigate the effects of numerical instability. The relative pose factor is defined:
Notably, in environments with well-distributed constraints, the pose prior carries minimal weight.
III-D2 Joint Optimization
We employ a factor graph approach to minimize all residual terms:
IV Experiment
In this section, we assess the viability of our proposed approach and evaluate its performance in real-world scenarios. Our analysis consists of three main stages:
Alignment Risks Prediction: We demonstrate the effectiveness of our alignment risk on potential failure scans including cave, tunnel, and long corridor environments. Additionally, we show the proposed confidence metric of the state’s direction to validate the consistency of our method in risk prediction.
Active Sensor Fusion: Building on the above covariance analysis, we demonstrate how to actively incorporate other prior odometry sources to overcome the degeneracy of cave, tunnel, and long corridor environments. We then compare the resulting map quality to show high accuracy and robustness.
Odometry Evaluation: Finally, we provide a thorough evaluation of the accuracy and robustness of our approach in challenging environments.
IV-A Alignment Risks Prediction
To evaluate the effectiveness of our approach, we conducted analyses across three challenging environments: caves, multi-floor structures, and long corridors (Fig. 1). For each setting, we aggregated approximately 10 seconds of point cloud data and projected our confidence metric onto the 3D space, creating an observability scan. In both cave and long-corridor environments, our platform’s forward movement revealed significantly lower confidence values in the X (forward) direction compared to other dimensions. This phenomenon is visually represented in the observability scans (Fig. 3), where the scarcity of red points indicates reduced correspondence in the X direction. The observability scans provide a clear visual representation of degeneracy in specific directions: a lack of red points (X direction) for caves and long corridors, and fewer blue points (Z direction) for multi-floor environments. This combination of quantitative low-confidence values and qualitative visual representation validates the consistency of our method in real-time alignment risk prediction.
IV-B Active Sensor Fusion
In this section, we demonstrate how to leverage confidence estimation in a timely way in order to incorporate supplementary odometry data to enhance localization robustness in these challenging settings. We conducted field experiments using handheld devices, legged robots, and ground vehicles. These tests took place in three degraded environments: i) narrow, self-similar cave environments, ii) confined, multi-floor environments, and iii) long corridors with flat open spaces. These scenarios pose significant challenges for LiDAR-based localization due to their repetitive geometries and lack of distinctive features. To assess the accuracy of our localization algorithm, we employed a FARO scanner to establish a ground truth map with a precision error of is less than 2mm. For each experimental run, we conducted map analysis to quantify the proportion of inliers and outliers in the localization-generated map relative to the ground truth map. We defined inliers and outliers using a 10 cm threshold. In each scenario, we compared our algorithm with HDL Localization [11] and FASTLIO Localization [12].
IV-B1 Cave Experiments
We evaluated our localization algorithm in Laurel Caverns, a challenging cave environment located in the suburbs of Pittsburgh, USA, as illustrated in Fig. 4. This site poses significant challenges for LiDAR SLAM systems due to its smooth, repetitive structures and poor lighting conditions. Despite these challenges, our proposed method demonstrated superior performance compared to prior approaches across multiple runs, as shown in Table I. For Cave01, it achieved an outlier rate of only 0.50% over 416 meters of travel, representing a 6-fold improvement over the second-best method (2.87% on the same run). This exceptional performance can be attributed to two key factors: i) effective prediction of scan registration risks before failures occur; and ii) accurate estimation of confidence in each state direction, enabling precise fusion of alternative odometry. As shown in Fig. 4, the confidence metric reveals that the LiDAR SLAM system’s performance degrades primarily in the X direction. Our method addresses this by selectively incorporating additional odometry data in this direction with appropriate confidence values.
IV-B2 Multi-Floor Environments
We evaluated our algorithm in multi-floor environments using a legged robot to navigate an abandoned hospital. This scenario presents unique challenges due to confined spaces like staircases, which restrict the sensor’s field of view. Moreover, even small errors in elevation estimation can lead to significant misalignments between floors, as shown in Fig. 1. Our method demonstrated superior performance in this challenging environment, Table I. It achieved an outlier rate of only 8.03% over 270 meters of travel, representing a 7.2-fold improvement over the second-best method (58.49%). This substantial enhancement in accuracy is particularly noteworthy given the complexity of the environment. Fig. 5 highlights a key feature of our approach: the blue dashed square at the bottom of the figure indicates an area of low confidence in the Z direction. In this region, our method actively incorporates visual odometry in the Z dimension.
IV-B3 Long Corridor Experiment
We used an RC car for the long corridor experiment, which operated at high speeds of up to 5 m/s and 57°/s, navigating both indoor corridors and outdoor open areas. Long corridors, with their repetitive and symmetrical features, such as identical walls and flat open spaces which lack distinct geometric features, present significant challenges for LiDAR systems (Fig. 6). Despite these challenges, our algorithm demonstrated a 7.8-fold improvement over the second-best method, Table I. This improvement is due not only to its ability to predict alignment risks but also to its capacity to identify the direction of degradation and assign appropriate confidence weights during sensor fusion. Notably, our method did not require any pre-defined heuristic thresholds to detect point cloud degeneration, even as the environment changed.
| % of Outliers | |||||||
| Method | Cave01 | Cave02 | Cave03 | Cave04 | Floor01 | Corridor01 | Corridor02 |
| Distance | 416m | 475m | 490m | 597m | 270m | 617m | 690m |
| HDL_LOC | 2.87 | 52.80 | 2.15 | 4.88 | 58.49 | 27.83 | 42.45 |
| FAST_LIO_LOC | 61.48 | 25.32 | 13.70 | 41.54 | 67.57 | 95.89 | 73.99 |
| SuperLoc | 0.50 | 0.81 | 1.50 | 1.60 | 8.03 | 3.55 | 23.99 |
| Method | Geometric Degradation (SubT-MRS) | Mix Degradation | Average | ||||||
| Final01 | Final02 | Final03 | Urban01 | Urban02 | Cave03 | Corridor01 | Floor01 | ||
| Liu* [33] | 0.307 | 0.095 | 0.629 | 0.122 | 0.235 | 0.260 | 1.454 | 0.401 | 0.588 |
| Weitong* [34] | 0.26 | 0.096 | 0.617 | 0.120 | 0.222 | 0.402 | 1.254 | 0.577 | 0.663 |
| Kim* [33] | 0.331 | 0.092 | 0.787 | 0.123 | 0.270 | 0.279 | 2.100 | 0.650 | 3.825 |
| Yibin [8] | 1.060 | 0.220 | 0.750 | 0.470 | 0.620 | 9.140 | 2.990 | 5.500 | 4.312 |
| Zhong* [35] | 1.205 | 0.695 | - | 1.175 | 1.72 | 2.08 | - | - | 1.209 |
| Zheng [18] | - | - | - | - | - | 3.786 | 55.205 | 19.769 | 26.254 |
| Our | 0.238 | 0.074 | 0.396 | 0.026 | 0.104 | 0.204 | 0.817 | 0.315 | 0.272 |
| Geometric Degradation (SubT-MRS) | Mix Degradation | |||||||||
| Method | Final01 | Final02 | Final03 | Urban01 | Urban02 | Cave03 | Corridor01 | Floor01 | Average | |
| Weitong*[34] | Rp | 0.922 | 0.929 | 0.906 | 0.933 | 0.919 | 0.830 | 0.889 | 0.909 | 0.905 |
| Liu*[33] | 0.922 | 0.929 | 0.906 | 0.933 | 0.919 | 0.830 | 0.885 | 0.905 | 0.904 | |
| Kim*[33] | 0.884 | 0.929 | 0.906 | 0.933 | 0.915 | 0.830 | 0.890 | 0.252 | 0.817 | |
| Yibin[8] | 0.849 | 0.897 | 0.827 | 0.875 | 0.795 | 0.751 | 0.502 | 0.738 | 0.779 | |
| Zhong* [35] | 0.278 | 0.910 | 0.827 | 0.905 | 0.877 | - | - | - | 0.759 | |
| Zheng[18] | - | - | - | - | - | 0.618 | 0.579 | 0.389 | 0.529 | |
| Our | 0.929 | 0.933 | 0.916 | 0.937 | 0.925 | 0.973 | 0.899 | 0.891 | 0.925 | |
| Weitong *[34] | Rr | 0.931 | 0.931 | 0.907 | 0.938 | 0.924 | 0.848 | 0.890 | 0.914 | 0.911 |
| Liu*[33] | 0.931 | 0.931 | 0.908 | 0.938 | 0.924 | 0.848 | 0.885 | 0.908 | 0.909 | |
| Kim*[33] | 0.923 | 0.931 | 0.908 | 0.938 | 0.920 | 0.848 | 0.890 | 0.260 | 0.827 | |
| Yibin[8] | 0.924 | 0.928 | 0.898 | 0.933 | 0.914 | 0.846 | 0.874 | 0.886 | 0.900 | |
| Zhong* [35] | 0.296 | 0.921 | 0.851 | 0.922 | 0.905 | - | - | - | 0.779 | |
| Zheng[18] | - | - | - | - | - | 0.629 | 0.882 | 0.489 | 0.667 | |
| Our | 0.938 | 0.942 | 0.925 | 0.943 | 0.941 | 0.964 | 0.944 | 0.924 | 0.940 | |
IV-C Odometry Accuracy Evaluation
To further validate the precision of pose estimation, we conducted Absolute Trajectory Error (ATE) analysis using our odometry system [15] on the SubT-MRS dataset [32]. This dataset encompasses challenging environments featuring sensor degradation, aggressive locomotion, and extreme weather conditions. The eight sequences in the dataset are categorized into two groups for testing: Geometric degradation (point cloud degradation only) and Mixed degradation (both point cloud and visual degradation). ATE results for competing systems were obtained from an open SLAM challenge [32]. Participating teams were permitted to employ loop closure techniques for post-processing accuracy enhancement. Despite this, our method still achieves superior consistency and accuracy, demonstrating an average ATE of 0.271 without employing loop closure (Table II). This performance surpasses the second-best method by 54%, underscoring our high accuracy.
IV-D Odometry Robustness Metric Evaluation
An ideal evaluation metric should consider both accuracy and completeness [32]. While ATE effectively assesses trajectory accuracy, it falls short in capturing completeness. To address this limitation, we employ robustness metrics[15] to consider both aspects. Tables III presents a comprehensive report of these metrics across various scenarios and algorithms. To see the value of these metrics, consider, e.g., that while the Floor01 ATE result for Zheng et al.’s method outperforms that of Corridor01, its value is worse due to the incomplete trajectory run. A summary of Robustness metrics across all other methods is shown in Fig. 7. The blue curve, representing the and of our odometry method, is the highest across almost the entire range.
IV-E Runtime Performance
Our proposed method achieves a processing rate of 22 frames per second (FPS), with each frame taking 45 ms to process on an AMD Ryzen 7 3700X 8-Core CPU.
V Conclusion
In this work, we proposed a map-based LiDAR localization that can anticipate and mitigate localization failures before the optimization process. Also, it provides a reliable observability estimation without requiring a heuristic threshold for different environments. Compared to existing methods, it achieves a remarkable 49.7% improvement in accuracy over the second-best algorithms. In future work, we plan to incorporate observability into global scan registration to achieve better re-localization for the kidnapping problem.